OpenStudy (anonymous):

how do i determine whether 27, 120, 123 are the measures of the sides of a right angle....and after that then state whether they form a Pythagorean triple?????......PLEASE HELP:)

OpenStudy (zarkon):

is \(27^2+120^2=123^2\)?

OpenStudy (zarkon):

also I assume you mean the sides of a right triangle

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

use the Pythagorean Theorem, like @Zarkon did

OpenStudy (anonymous):

o okay thank you both

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\] where "c" is the longest side (aka the hypotenuse) and "a" & "b" are the other side lengths (aka the legs)

OpenStudy (anonymous):

So: 27, 120, 123 \[27^2 + 120^2 = 123^2\] \[729 + 14400 = 15129\] \[15129 = 15129\] THEREFORE, the side lengths given (27, 120, 123) can form a right triangle.